News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 21 June 2002 8.00am EDT

During the last week, Chandra completed the observing schedule as planned.

A real-time procedure was completed successfully on June 17 to check the bright Earth limb limitation with the aspect camera during perigee passage. One issue identified during the procedure was that a series of hardware commands sent from the daily load to the aspect camera to acquire monitor windows was not executed. The issue is under investigation.

Initial assessment of the results from the Earth limb check shows consistency with prior measurements made in November 2001. The camera remains locked on stars to within ~4-5 degrees of the bright earth (this compares with the 20 degree current limitation). Further measurements are planned before a constraint change can be considered.

The observing schedule for the next week is shown below and includes two observations of Her X-1 that are phase constrained.

BRI2235-0301 Jun 23 1AXG J010952-1252 ACIS-I AX J0146.2-0354 ACIS-S Jun 24 A133 ACIS-I NGC 985 ACIS-S/HETG Radiation Belts Jun 25 RXJ0027.6+2616 ACIS-I Jun 26 1AXG J003331-2139 ACIS-I HS 2135+1326 ACIS-S A133 ACIS-I HD 150136 ACIS-S/HETG Jun 27 Radiation Belts Jun 28 GRS 1758-258 HRC-I Her X-1 ACIS-S/HETG AX J1354.0+3346 ACIS-S Jun 29 Her X-1 HRC-S/LETG MACSJ1427.6-2521 ACIS-I TON 1388 HRC-S/LETG IRAS 11119+3257 ACIS-S Jun 30


All spacecraft subsystems continued to operate nominally.

Return to Status Reports